Minutes ITBApril 282011

Introduction

  • Attendees: Suchandra, Anupam, Xin, Steve,
  • Apologies: Scot, Alain, Horst, Karthik
  • Coordinates: Thursday, 2:30pm Central; 1-866-740-1260 Meeting ID 4503064

Persistent ITB Sites (Admins)

Last Update

  • BNL
  • FNAL
    • Doing miscellaneous testing with new certs
    • No other updates
  • OUHEP
    • No updates
    • Updated to condor 7.6.0, the change appears to have fixed htpc configuration
  • UC_ITB
    • Machines still being brought up from move
    • Should be up this afternoon
  • VDT
    • No change

This Week

  • BNL
    • No changes in software
    • Had problems with the ITB gatekeeper in the last 2 days or so * prima callout failing when talking to GUMS * CE not accessible
  • FNAL
    • No changes recently
  • OUHEP
  • UC_ITB
    • No changes
    • Hadoop/bestman2 SE being brought online
  • VDT
    • No updates

VTB/ITB testing

  • Worker node rpms
    • FNAL - already testing
    • BNL, UC - will test
  • CMS requested changes
    • sleep added to fork jobmanager
    • globus-gatekeeper will be niced more to give it lower priority
    • patch to support group accounts in the condor jobmanager
    • scripts to cleanup globus files left by jobs

HTPC

  • Site status:
    • OUHEP
      • Updated to condor 7.6.0
    • BNL
    • UC_ITB_CONDOR
      • Cluster is still down
    • FNAL
      • Will test HTPC but not in near term (about a month from here)

Documentation

AOB

  • None

-- SuchandraThapa - 27 Apr 2011

Topic revision: r2 - 28 Apr 2011 - 19:48:59 - SuchandraThapa

Hello, TWikiGuest!
Register

 
TWIKI.NET

TWiki | Report Bugs | Privacy Policy

This site is powered by the TWiki collaboration platformCopyright by the contributing authors. All material on this collaboration platform is the property of the contributing authors..